A laser of wavelength 532 nm is shot through a double slit at a screen 4.4 m away. If the separation between slits is 0.69 mm, what is the spacing between the fringes resultant on the screen?
1
Expert's answer
2019-09-23T09:17:59-0400

The distance from the center of the interference picture to the mth maximum is "y = \frac{m \lambda D}{d}", where "D" is the distance from the slit to the screen, "d" is the distance between two slits, "\lambda" is the wavelength. The distance between the fringes is the distance between two neighboring maxima/minima. Hence, according to the formula, mentioned above, the distance is "\Delta = \frac{\lambda D}{d} = 3.4 \cdot 10^{-3} m = 3.4 mm".